Water flowing along a horizontal pipe has a speed of 2.27 m/s and a pressure of 4.86 atm. Further along, the pipe narrows so that the cross-sectional area decreases by a factor of 6.44. What is the pressure in the narrow section?


1
Expert's answer
2021-01-21T09:57:12-0500

"V_{1}=2.27m\/s, P_{1}=4.86 atm"

cross-sectional area decreased by a factor of 6.44

"A_{1}\/A_{2}=6.44"

"A_{1}V_{1}=A_{2}V_{2}"

"V_{2}=(A_{1}V_{1})/A_{2}=6.44\times2.27=14.6188m/s"

From Bernoulli's theorem,

"\Delta P=\rho (V_{1}^{2}-V_{2}^{2})/2"

"P_2-P_1 =1000\times(2.27^{2}-14.6188^{2})/2"

"P_2-P_1=-104278.2067P_a=-\frac{104278.2067}{101325}atm=-1.029atm"

"P_2=4.86-1.029=3.831atm"